Barracuda Inc., has a beta of 1.40, the annual risk free rate of interest is currently 10
percent, and the required return on the market portfolio is 16 percent. The firm
estimates that its future dividends will continue to increase at an annual compound rate
consistent with that experienced over the 2009–2012 period.
Year Dividend(RM)
2009 2.70
2010 2.95
2011 3.25
2012 3.40
(a) Estimate the value of Barracuda Inc., stock.

(b) A lawsuit has been filed against the company by a competitor, and the potential loss
has increased risk, which is reflected in the company’s beta, increasing it to 1.6. What
is the estimated price of the stock following the filing of the lawsuit.
1
Expert's answer
2016-05-11T08:37:02-0400
Beta b = 1.40, rf = 10 percent, the required return on the market portfolio is 16 percent.
So, as r = rf + b*(rm - rf), then:
0.16 = 0.1 + 1.4*(rm - 0.1),
rm = 0.06/1.4 + 0.1 = 0.143 = 14.3%.
Year Dividend(RM)
2009 2.70
2010 2.95
2011 3.25
2012 3.40
(a) To calculate the value of Barracuda Inc. stock, firstly we should calculate the average growth rate g = (2.95/2.7 + 3.25/2.95 + 3.4/3.25)/3 - 1 = 0.08 = 8%.
Price of stock is P = Do*(1 + g)/(r - g) = 2.7*1.08/(0.16 - 0.08) = $36.45.
(b) If a lawsuit has been filed against the company by a competitor, and the potential loss has increased risk, which is reflected in the company’s beta, increasing it to 1.6, then:
New required rate of return is:
r = 0.1 + 1.6*(0.143 - 0.1) = 0.1688 = 16.88%.
New price of stock is P = Do*(1 + g)/(r - g) = 2.7*1.08/(0.1688 - 0.08) = $32.84,
so, the estimated price of the stock following the filing of the lawsuit will decrease.